OpenStudy (anonymous):

The equation of a curve C is y=2x^2−8x+9 and the equation of a line L is x+y=3. Find the x-coordinates of the points of intersection of L and C.

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

y = 2x^2-8x+9 x+y=3 y = y so 2x^2-8x+9=3-x 2x^2-7x+6=0 find the roots and be happy

OpenStudy (anonymous):

Oh, I got them. Thanks

OpenStudy (anonymous):

welcome